OpenStudy (anonymous):

For what values of x is f (x)=0 when f(x)= x^2-x-6?

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{f(x)= x^2-x-6\\~\\ f (x)=0 ----(given)\\~\\ therefore\\~\\ x^2-x-6=0\\~\\ x^2-3x+2x-6=0\\~\\ x(x-3)+2(x-3)=0\\~\\ (x-3)(x+2)=0\\~\\ (x-3)=0~~or~~ (x+2)=0\\~\\ x=3~~or~~ x=-2\\~\\~\\ }\end{align}\)